### v1

@@ Abstract
With the ever-increasing number of news stories available online, classifying them by topic, regardless of the language they are written in, has become crucial for enhancing readers’ access to relevant content. To address this challenge, we propose a teacher-student framework based on large language models (LLMs) for developing multilingual news classification models of reasonable size with no need for manual data annotation. The framework employs a Generative Pretrained Transformer (GPT) model as the teacher model to develop an IPTC Media Topic training dataset through automatic annotation of news articles in Slovenian, Croatian, Greek, and Catalan. The teacher model exhibits a high zero-shot performance on all four languages. Its agreement with human annotators is comparable to that between the human annotators themselves. To mitigate the computational limitations associated with the requirement of processing millions of texts daily, smaller BERT-like student models are fine-tuned on the GPT-annotated dataset. These student models achieve high performance comparable to the teacher model. Furthermore, we explore the impact of the training data size on the performance of the student models and investigate their monolingual, multilingual and zero-shot cross-lingual capabilities. The findings indicate that student models can achieve high performance with a relatively small number of training instances, and demonstrate strong zero-shot cross-lingual abilities. Finally, we publish the best-performing news topic classifier, enabling multilingual classification with the top-level categories of the IPTC Media Topic schema.
@@ I Introduction
Topic classification is a significant asset in the news industry, enabling the automatic identification of news topics and facilitating readers’ access to content that aligns with their interests. However, developing a robust news topic classifier is challenging, particularly because of the absence of manually-annotated data, which are especially scarce for non-English languages. Manual annotation, although invaluable, is an expensive and labor-intensive process that cannot be quickly extended to numerous languages.
@@
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=0.9]kuzma1.png Pipeline for the preparation of training and test datasets for the news topic classification.
Despite the emergence of zero-shot approaches employing Generative Pretrained Transformers (GPTs) that have demonstrated impressive results in various natural language processing tasks in diverse languages [1], [2, 3], their application in settings with extensive data to be processed remains impractical because of their high computational demands. A more suitable technology for this setting is a smaller BERT-like language model; however, such models require thousands of annotated instances for effective fine-tuning for a certain task.
@@
In our experiments, we train and evaluate the models on multilingual training and test datasets comprising four diverse languages: Catalan, Croatian, Greek, and Slovenian. The selection of these languages for evaluation is on one side based on their availability in the MaCoCu corpora collection [5] to ensure a high level of comparability between language-specific datasets. Furthermore, these languages exhibit varying degrees of linguistic relatedness, with Croatian and Slovenian being closely related and the other two languages belonging to separate branches of the Indo-European language family. This distinction provides valuable insights for cross-lingual experiments.
In this work, we aim to address the following research questions, pertaining to the task of IPTC news topic classification:
@@
    How much data, annotated by the teacher model, is required for the student model to achieve comparable performance to the teacher model?
@@
In summary, this study makes the following contributions. We investigate the feasibility of the LLM teacher-student framework for the development of accurate and computationally efficient multilingual news topic classifiers for languages that do not have readily accessible manually-annotated training data suitable for the task. Specifically, we focus on single-label multi-class text classification employing top-level categories of the IPTC NewsCodes Media Topic schema. This study assesses the applicability of a GPT model for data annotation by comparing the level of agreement between the model and human annotators to the level of agreement among the human annotators themselves. Furthermore, we examine the classification performance of the student model relative to its teacher model, while also exploring the impact of training data size on the student model’s performance, as well as its zero-shot cross-lingual and multilingual capabilities. Motivated by these positive results, we publish the best-performing model on the Hugging Face repository (https://huggingface.co/classla/multilingual-IPTC-news-topic-classifier), providing an IPTC news topic classification model that can be applied to any of the 100 languages covered by the XLM-RoBERTa model. To the best of our knowledge, this is the first openly-available multilingual topic classifier that uses the (top-level) IPTC Media Topic categories.
The remainder of this paper is organized as follows. Section II provides an overview of the existing literature on IPTC news topic classification and the application of GPT models to data annotation. In Section III, we introduce our methodology for the automatic annotation of training data by leveraging the zero-shot capabilities of large language models. In Section IV, we describe the manual annotation campaign that was undertaken to develop a test set for evaluating the performance of both teacher and student models, and we evaluate the performance of the GPT model as a data annotator in comparison to human annotators. Section V presents the experiments involving fine-tuning BERT-like models on varying sizes of training data, and the analysis of their monolingual, multilingual, and zero-shot cross-lingual performance. Section VI concludes the paper with a discussion of the main findings and suggestions for future research. Additionally, the Appendix provides more details on the annotation guidelines, and the description of the top-level IPTC Media Topic labels (Section A), and the prompt used for automatic data annotation (Section B).
@@ II Related Work
The International Press Telecommunications Council (IPTC) is a global organization dedicated to the development and promotion of industry standards for the exchange of news data. Among its notable contributions is the maintenance of IPTC NewsCodes controlled vocabularies that are widely adopted by major media providers, including Agence France-Presse, Associated Press, and Reuters [6]. The NewsCodes vocabularies include the Media Topic vocabulary, which has been established as a global standard to ensure consistent coding of news metadata across diverse news providers [7]. The Media Topic taxonomy is structured hierarchically, with 17 topic labels at the top level. The entire taxonomy encompasses over 1,000 topic labels, organized across up to five levels of granularity [8].
Multiple studies researched news topic classification [9, 10], [11, 12], [13]. However, there is no manually-annotated reference dataset that can be used directly for the training and evaluation of IPTC Media Topic classifiers. The existing topic datasets have several limitations:
@@
    Inconsistent Topic Schemata: Many datasets use their own topic schemata [12, 9, 13], rendering the results of their experiments highly dataset-dependent, and consequently, incomparable with other studies.
@@
Recently, two news topic datasets have been introduced, both annotated with the IPTC Media Topic schema: the MN-DS dataset [17], comprising approximately 10,000 English news articles, and the EventDNA dataset [18], containing around 1,800 Dutch news articles. Despite their potential, both datasets have faced criticism regarding the reliability of manually-annotated labels, as highlighted in related studies employing these datasets for machine learning experiments [19, 20]. Beyond problems with reliability, these datasets exhibit additional limitations. The MN-DS dataset was derived from the NELA-GT-2018 dataset [21], which was originally developed for misinformation research. As a result, it includes sources known to disseminate fake news, which may compromise the performance of models trained on it when applied to mainstream news texts. The EventDNA dataset, on the other hand, is limited by its annotations being performed at the lowest levels of the IPTC Media Topic taxonomy. Specific sublabels do not always align well with top-level labels, which restricts the utility of the dataset for automatic text classification on top-level labels. Additionally, while EventDNA is freely accessible, restrictions on third-party sharing pose challenges for conducting experiments with closed-source large language models.
@@
Recently introduced Generative Pretrained Transformer (GPT) models, commonly known as large language models (LLMs), have demonstrated impressive performance across a range of text classification tasks, even when used in a zero-shot prompting fashion that does not require any training data [1, 22, 23]. Furthermore, recent research in the field of natural language processing (NLP) has embraced “LLMs as catalysts for redefining the landscape of data annotation in machine learning and NLP” [24]. Specifically, researchers are exploring the potential of GPT-based annotation of training data to replace time-consuming manual annotation, with the aim of using these annotations to fine-tune or improve the performance of other GPT or BERT-like models [24, 25], [26, 27]. Although LLMs have been widely adopted to generate and annotate instruction-tuning datasets [28], [29, 30], their use in the annotation of text classification data has been less explored. Meng et al. [31] contributed to this area by employing a GPT model to generate both texts and labels for training datasets through a zero-shot prompting approach, and by fine-tuning a smaller language model on the generated training dataset. In contrast to this approach, our study refrains from generating synthetic data to mitigate the potential LLM biases. Instead, we use authentic news articles sourced from the web and employ a GPT model exclusively for classifying these texts into Media Topic labels. This GPT-based data annotation approach was also shown to be promising in the domain of sentiment analysis, where student models fine-tuned on datasets annotated by a GPT model exhibited comparable performance to models fine-tuned on datasets annotated by human annotators [27].
@@
A recent investigation into the applicability of GPT models for IPTC news topic classification was conducted by Fatemi et al. [20] who evaluated a GPT model on the MN-DS dataset [17]. The applicability of GPT models for topic annotation was evaluated by training various machine learning models on the GPT-annotated dataset and comparing their performance against models trained on the same dataset, but manually-annotated. The findings indicate the promising performance of the GPT model used as a data annotator. However, it is important to note that the performance of the model was not directly evaluated on a manually-annotated test set. Consequently, the accuracy of the GPT model in this task remains undetermined.
@@
Building on existing research, our study explores the use of GPT-based data annotation to fine-tune a BERT-like topic classifier. This research addresses multiple gaps in the literature. First, we assess the performance of both the teacher GPT model and fine-tuned student models on reliably manually-annotated data. Second, we compare the annotation reliability of the GPT model with the reliability of human annotators. Additionally, our research extends machine learning experiments across multiple languages, offering valuable insights into the models’ multilingual and cross-lingual performance.
@@ III LLM-Based Dataset Development
In this section, we present our methodology for the automatic annotation of training data aimed at fine-tuning a multilingual Transformer model that employs top-level IPTC Media Topic labels. The proposed methodology leverages three resources that have become available recently:
@@
  • Multilingual GPT Models: The multilingual Generative Pretrained Transformer (GPT) models have demonstrated impressive zero-shot classification capabilities across numerous languages and tasks. Inter alia, these models have shown high performance on South Slavic languages, including Croatian and Slovenian that are included in our experiments [34]. Specifically, we employ the GPT-4o model (version gpt-4o-2024-05-13) [35] provided by the OpenAI API, which outperformed other open- and closed-source GPT models in our preliminary experiments.
@@ III-A Extraction of News Texts and Pre-Processing
The pipeline for developing training and test data for news topic classification using the LLM teacher-student framework is illustrated in Fig. I. The datasets used in this study are sourced from the Catalan (ca) [36], Croatian (hr) [37], Greek (el) [38], and Slovenian (sl) [39] MaCoCu web corpora that have been annotated with genres (genre-annotated datasets are available as part of the MaCoCu-Genre corpus collection [40]). Given the focus of the IPTC topic classification on news articles, we extract from the web corpora the subset of news articles, identified with the X-GENRE classifier [1]. Furthermore, the dataset undergoes additional pre-processing, wherein the texts are truncated to the initial 512 words, which is a constraint imposed by the BERT-like models.
@@ III-C Training and Development Datasets
The distribution of labels within the EMMediaTopic dataset is shown in Fig. III. The dataset is shown to be relatively balanced by labels, with the most prevalent label (sport) accounting for approximately 15% of the instances. Moreover, the distribution of labels is consistent across the four languages. The least represented labels are weather, accounting for less than 1% of instances, and religion and labour, each assigned to less than 2% of texts.
@@ III-D Test Dataset
The test dataset is derived from the second batch of 8,000 automatically annotated texts, comprising texts in the same four languages as the training data, namely Slovenian, Croatian, Catalan, and Greek. To assess the models’ performance across all 17 top IPTC Media Topic labels, the test set is balanced by GPT-assigned labels which is achieved by selecting 18 instances per label per language (or fewer if there are fewer than 18 label instances in the batch). The selection of texts to be manually-annotated comprised 1,199 instances with a balanced distribution across languages and GPT-assigned labels. Finally, the test set was manually annotated to obtain human gold labels. Details on the annotation campaign are provided in the next section.
@@ III-E Teacher Model Performance
The first application of the manually-annotated test set was to measure the performance of the GPT-4o model that was used for automatic annotation. The model demonstrates a consistently high performance, with an average macro-F1 score of 0.731 and a micro-F1 score of 0.722. Its performance across different languages is stable, with a 5-point difference between the highest macro-F1 score (on Slovenian) and the lowest (on Catalan), as shown in Table I.
@@ IV Manual Annotation of Test Data
The test sample, balanced by GPT-assigned labels and the four languages, is manually annotated by a single annotator. The annotator was provided with annotation guidelines with the description of the labels (see Appendix A) and received brief training on the annotation process. The description of the labels is based on the official definitions provided by the IPTC [42]. However, as these definitions offer only a high-level understanding of the distinctions between labels, we enriched the descriptions with examples of their lower-level labels to facilitate a more precise and consistent annotation.
In addition to the standard 17 top-level IPTC Media Topic labels, three additional labels have been introduced to assist the annotator in identifying and excluding texts that are unsuitable for the purposes of our task: do not know (highly challenging instance lacking a clear topic), not news (text of a different genre, e.g., Promotion), and multiple (presence of multiple texts in a single instance). The manual annotation process excluded 70 instances (5.83%) as unsuitable. Among these, 4% were excluded because of the absence of a clear topic, 2% because they contained multiple texts in a single instance, and 0.1% because of incorrect genre classification. The final test sample consists of 1,129 text instances. The test dataset is available on request from the corresponding author. It is not publicly released to prevent its integration into large language models (LLMs) during their pretraining or fine-tuning phase to maintain the integrity of future model evaluations.

### v2

@@ Abstract
With the ever-increasing number of news stories available online, classifying them by topic, regardless of the language they are written in, has become crucial for enhancing readers’ access to relevant content. To address this challenge, we propose a teacher-student framework based on large language models (LLMs) for developing multilingual news topic classification models of reasonable size with no need for manual data annotation. The framework employs a Generative Pretrained Transformer (GPT) model as the teacher model to develop a news topic training dataset through automatic annotation of 20,000 news articles in Slovenian, Croatian, Greek, and Catalan. Articles are classified into 17 main categories from the Media Topic schema, developed by the International Press Telecommunications Council (IPTC). The teacher model exhibits high zero-shot performance in all four languages. Its agreement with human annotators is comparable to that between the human annotators themselves. To mitigate the computational limitations associated with the requirement of processing millions of texts daily, smaller BERT-like student models are fine-tuned on the GPT-annotated dataset. These student models achieve high performance comparable to the teacher model. Furthermore, we explore the impact of the training data size on the performance of the student models and investigate their monolingual, multilingual, and zero-shot cross-lingual capabilities. The findings indicate that student models can achieve high performance with a relatively small number of training instances, and demonstrate strong zero-shot cross-lingual abilities. Finally, we publish the best-performing news topic classifier, enabling multilingual classification with the top-level categories of the IPTC Media Topic schema.
@@ I Introduction
Topic classification is a significant asset in the news industry, enabling automatic identification of news topics and facilitating readers’ access to content that aligns with their interests. However, developing a robust news topic classifier is challenging, particularly because of the absence of manually annotated data, which are especially scarce for non-English languages. Manual annotation, although invaluable, is an expensive and labor-intensive process that cannot be quickly extended to numerous languages.
@@
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=0.9]kuzma1.png Pipeline for the preparation of training and test datasets for the classification of news topics.
Despite the emergence of zero-shot approaches that employ Generative Pretrained Transformers (GPTs) that have demonstrated impressive results in various natural language processing tasks in diverse languages [1], [2, 3], their application in settings with extensive data to be processed remains impractical due to their high computational demands. A more suitable technology for this setting is a smaller BERT-like language model; however, such models require thousands of annotated instances for effective fine-tuning for a certain task.
@@
In our experiments, we train and evaluate the models on multilingual training and test datasets comprising four diverse languages: Catalan, Croatian, Greek, and Slovenian. The selection of these languages for evaluation is based on their availability in the MaCoCu corpora collection [5] to ensure a high level of comparability between language-specific datasets. Furthermore, these languages exhibit varying degrees of linguistic relatedness, with Croatian and Slovenian being closely related, and the other two languages belonging to separate branches of the Indo-European language family. This distinction provides valuable insights for cross-lingual experiments.
In this work, our aim is to address the following research questions, pertaining to the task of IPTC news topic classification:
@@
    How much data, annotated by the teacher model, are required for the student model to achieve performance comparable to the teacher model?
@@
In summary, this study makes the following contributions. We investigate the feasibility of the LLM teacher-student framework for the development of accurate and computationally efficient multilingual news topic classifiers for languages that do not have readily accessible manually annotated training data suitable for the task. Specifically, we focus on single-label multi-class text classification employing top-level categories of the IPTC NewsCodes Media Topic schema. This study assesses the applicability of a GPT model for data annotation by comparing the level of agreement between the model and human annotators to the level of agreement among the human annotators themselves. Furthermore, we examine the classification performance of the student model relative to its teacher model, while also exploring the impact of training data size on the student model’s performance, as well as its zero-shot cross-lingual and multilingual capabilities. Motivated by these positive results, we publish the best performing model in the Hugging Face repository (https://huggingface.co/classla/multilingual-IPTC-news-topic-classifier), providing an IPTC news topic classification model that can be applied to any of the 100 languages covered by the XLM-RoBERTa model. To the best of our knowledge, this is the first openly-available multilingual topic classifier that uses the (top-level) IPTC Media Topic categories.
The remainder of this paper is organized as follows. Section II provides an overview of the existing literature on IPTC news topic classification and the application of GPT models to data annotation. In Section III, we introduce our methodology for automatic annotation of training data by leveraging the zero-shot capabilities of large language models. In Section IV, we describe the manual annotation campaign that was undertaken to develop a test set for evaluating the performance of both teacher and student models, and we evaluate the performance of the GPT model as a data annotator compared to human annotators. Section V presents the experiments that involve fine-tuning BERT-like models on varying sizes of training data, and the analysis of their monolingual, multilingual, and zero-shot cross-lingual performance. Section VI concludes the paper with a discussion of the main findings and suggestions for future research. Additionally, the Appendix provides more details on the annotation guidelines and the description of the top-level IPTC Media Topic labels (Section A), and the prompt used for automatic data annotation (Section B).
@@ II Related Work
The International Press Telecommunications Council (IPTC) is a global organization dedicated to the development and promotion of industry standards for the exchange of news data. Among its notable contributions is the maintenance of IPTC NewsCodes controlled vocabularies that are widely adopted by major media providers, including Agence France-Presse, Associated Press, and Reuters [6]. The NewsCodes vocabularies include the Media Topic vocabulary, which has been established as a global standard to ensure consistent coding of news metadata across diverse news providers [7]. The Media Topic taxonomy is structured hierarchically, with 17 topic labels at the top level. The entire taxonomy encompasses more than 1,000 topic labels, organized across up to five levels of granularity [8].
Multiple previous studies investigated news topic classification [9, 10], [11, 12], [13]. However, there is no manually annotated reference dataset that can be used directly for the training and evaluation of IPTC Media Topic classifiers. The existing topic datasets have several limitations:
@@
    Inconsistent Topic Schemata: Many datasets use their own topic schemata [12, 9, 13], rendering the results of their experiments highly dataset dependent and consequently incomparable with other studies.
@@
Recently, two news topic datasets have been introduced, both annotated with the IPTC Media Topic schema: the MN-DS dataset [17], comprising approximately 10,000 English news articles, and the EventDNA dataset [18], containing around 1,800 Dutch news articles. Despite their potential, both datasets have faced criticism regarding the reliability of manually annotated labels, as highlighted in related studies that employed these datasets for machine learning experiments [19, 20]. Beyond problems with reliability, these datasets exhibit additional limitations. The MN-DS dataset was derived from the NELA-GT-2018 dataset [21], which was originally developed for misinformation research. As a result, it includes sources known to disseminate fake news, which may compromise the performance of models trained on it when applied to mainstream news texts. The EventDNA dataset, on the other hand, is limited by its annotations being performed at the lowest levels of the IPTC Media Topic taxonomy. Specific sublabels do not always align well with top-level labels, which makes the dataset less useful for automatic text classification on top-level labels. Additionally, while EventDNA is freely accessible, restrictions on third-party sharing do not allow experiments with closed-source large language models.
@@
Recently introduced Generative Pretrained Transformer (GPT) models, commonly known as large language models (LLMs), have demonstrated impressive performance across a range of text classification tasks, even when used in a zero-shot prompting fashion that does not require any training data [1, 22, 23]. Furthermore, recent research in the field of natural language processing (NLP) has embraced “LLMs as catalysts for redefining the landscape of data annotation in machine learning and NLP” [24]. Specifically, researchers are exploring the potential of GPT-based annotation of training data to replace time-consuming manual annotation, with the aim of using these annotations to fine-tune or improve the performance of other GPT or BERT-like models [24, 25], [26, 27]. Although LLMs have been employed to generate and annotate instruction-tuning datasets [28], [29, 30], their use for the annotation of text classification data has been less explored. Meng et al. [31] contributed to this area by employing a GPT model to generate both texts and labels for training datasets through a zero-shot prompting approach, and by fine-tuning a smaller language model on the generated training dataset. In contrast to this approach, our study refrains from generating synthetic data to mitigate potential LLM biases. Instead, we use authentic news articles sourced from the web and employ a GPT model exclusively for classifying these texts into Media Topic labels. This GPT-based data annotation approach was also shown to be promising in the domain of sentiment analysis, where student models fine-tuned on datasets annotated by a GPT model exhibited comparable performance to models fine-tuned on datasets annotated by human annotators [27].
@@
A recent investigation into the applicability of GPT models for IPTC news topic classification was conducted by Fatemi et al. [20] who evaluated a GPT model on the MN-DS dataset [17]. The applicability of GPT models for topic annotation was evaluated by training various machine learning models on the GPT-annotated dataset and comparing their performance against models trained on the same dataset, but manually annotated. The findings indicate the promising performance of the GPT model used as a data annotator. However, it is important to note that the performance of the model was not directly evaluated on a manually annotated test set. Consequently, the accuracy of the GPT model in this task remains undetermined.
@@
Building on existing research, our study explores the use of GPT-based data annotation to fine-tune a BERT-like topic classifier. This research addresses multiple gaps in the literature. First, we assess the performance of both the teacher GPT model and fine-tuned student models on reliably manually annotated data. Second, we compare the annotation reliability of the GPT model with the reliability of human annotators. Additionally, our research involves machine learning experiments on multiple languages, which offer valuable insights into the models’ multilingual and cross-lingual performance.
@@ III LLM-Based Dataset Development
In this section, we present our methodology for the automatic annotation of training data that is used for fine-tuning a multilingual Transformer model which employs top-level IPTC Media Topic labels. The proposed methodology leverages three resources that have recently become available:
@@
  • Multilingual GPT Models: The multilingual Generative Pretrained Transformer (GPT) models have demonstrated impressive zero-shot classification capabilities across numerous languages and tasks. Inter alia, these models have shown high performance on South Slavic languages, including Croatian and Slovenian, which are included in our experiments [34]. Specifically, we employ the GPT-4o model (version gpt-4o-2024-05-13) [35] provided by the OpenAI API, which outperformed other open- and closed-source GPT models in our preliminary experiments.
@@ III-A Extraction of News Texts and Pre-Processing
The pipeline for developing training and test data for news topic classification using the LLM teacher-student framework is illustrated in Fig. I. The datasets used in this study are sourced from the Catalan (ca) [36], Croatian (hr) [37], Greek (el) [38], and Slovenian (sl) [39] MaCoCu web corpora that have been annotated with genres (genre-annotated datasets are available as part of the MaCoCu-Genre corpus collection [40]). Given the focus of IPTC topic classification on news articles, we extract from the web corpora the subset of news articles, identified with the X-GENRE classifier [1]. The X-GENRE classifier provides highly reliable classification into the News genre, as well as eight other genre categories. Evaluation on a multilingual, manually annotated test set, sampled from the MaCoCu corpora, revealed F1 scores for the News label ranging from 0.82 in Catalan to 0.95 in Croatian [41]. Furthermore, the dataset undergoes additional pre-processing, wherein the texts are truncated to the initial 512 words, which is a constraint imposed by the BERT-like models.
@@ III-C Training and Development Datasets
The distribution of labels within the EMMediaTopic dataset is shown in Fig. III. The dataset is shown to be relatively balanced by labels, with the most prevalent label (sport) accounting for approximately 15% of the instances. Moreover, the distribution of labels is consistent across the four languages. The least represented labels are weather, accounting for less than 1% of instances, and religion and labour, each assigned to less than 2% of the texts.
@@ III-D Test Dataset
The test dataset is derived from the second batch of 8,000 automatically annotated texts, comprising texts in the same four languages as the training data, namely Slovenian, Croatian, Catalan, and Greek. To assess the models’ performance across all 17 top IPTC Media Topic labels, the test set is balanced by GPT-assigned labels, which is achieved by selecting 18 instances per label per language (or fewer if there are fewer than 18 label instances in the batch). The selection of texts to be manually annotated comprised 1,199 instances with a balanced distribution across languages. Finally, the test set was manually annotated to obtain human gold labels. Details on the annotation campaign are provided in the next section.
@@ III-E Teacher Model Performance
The first application of the manually annotated test set was to measure the performance of the GPT-4o model that was used for automatic annotation. The model demonstrates consistently high performance, with an average macro-F1 score of 0.731 and a micro-F1 score of 0.722. Its performance across different languages is stable, with a 5-point difference between the highest macro-F1 score (on Slovenian) and the lowest (on Catalan), as shown in Table I.
@@ IV Manual Annotation of Test Data
The test sample, balanced by GPT-assigned labels and the four languages, is manually annotated by a single annotator. The annotator was provided with annotation guidelines with the descriptions of the labels (see Appendix A) and received brief training on the annotation process. The description of the labels is based on the official definitions provided by the IPTC [43]. However, since these definitions offer only a high-level understanding of the distinctions between labels, we enriched the descriptions with examples of their lower-level labels to facilitate a more precise and consistent annotation.
In addition to the standard 17 top-level IPTC Media Topic labels, three additional labels have been introduced to assist the annotator in identifying and excluding texts that are not suitable for the purposes of our task: do not know (highly challenging instance lacking a clear topic), not news (text of a different genre, e.g., Promotion), and multiple (presence of multiple texts in a single instance). The manual annotation process excluded 70 instances (5.83%) as unsuitable. Among these, 4% were excluded because of the absence of a clear topic, 2% because they contained multiple texts in a single instance, and 0.1% due to incorrect genre classification. The final test sample consists of 1,129 text instances. The test dataset is available on request from the corresponding author. It is not publicly released to prevent its integration into large language models (LLMs) during their pretraining or fine-tuning phase to maintain the integrity of future model evaluations.